OpenStudy (awesome781):

Two snowballs with masses of 0.40 kg and 0.60 kg, respectively, collide head-on and combine to form a single snowball. The initial speed for each is 15 m/s. If the velocity of the snowball with a mass of 1.0 kg is 3.0 m/s after the collision, what is the decrease in kinetic energy?

OpenStudy (fifciol):

Kinetic energy before: \[KE_b=\frac{ 1 }{2 }m_1v_1^2+\frac{ 1 }{2 }m_2v_2^2\] Kinetic energy after: \[KE_a=\frac{ 1 }{2 }m_3v_3^2\] decrease in KE: \[\Delta KE=KE_b-KE_a \] m1=0.4kg m2= 0.6kg m3=1kg v1=v2=15m/s v3=3m/s

OpenStudy (awesome781):

So the answer is 3.0 m/s?

OpenStudy (fifciol):

No ,the answer must be in joules. KE before the collision is simply the sum of kinetic energies of each ball

OpenStudy (fifciol):

this is 0.5*0.4*225 +0.5*0.6*225

OpenStudy (awesome781):

Ok I still don't understand....

OpenStudy (fifciol):

the kinetic energy before collision is then 112.5 J

OpenStudy (fifciol):

after the collision you no longer have two balls but one with mass 1kg and speed 3m/s

OpenStudy (fifciol):

the kinetic energy of that system is 0.5*1*9=4.5J

OpenStudy (awesome781):

The answer key that I have says 110 j

OpenStudy (fifciol):

so you've lost 112.5-4.5 = 108 J

OpenStudy (fifciol):

could be rounded off to 110 J

OpenStudy (awesome781):

Ok

OpenStudy (fifciol):

yw:)